ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 189, 2012

Aljadeff, E. Group graded P I-algebras and their codimension growth

189

Balister, P. and R. McCutcheon A concentration function estimate and intersective sets from matrices

413

Bollobás, B., I. Leader and M. Walters Lion and man—can both win?

267

Brehm, U. and W. Kühnel Lattice triangulations of 𝔼^{3} and of the 3-torus

97

Cardin, P. T., P. R. da Silva and M. A. Teixeira Implicit differential equations with impasse singularities and singular perturbation problems

307

Downarowicz, T. and Y. Lacroix Topological entropy zero and asymptotic pairs

323

Gray, R. and N. Ruskuc On maximal subgroups of free idempotent generated semigroups

147

Hamann, M. End-transitive graphs  437

Hegyvári, N. and F. Hennecart A note on Freiman models in Heisenberg groups

397

Hennecart, F. See N. Hegyvári Isaacs, I. M., G. Navarro and J. Sangroniz p-groups having few

almost-rational irreducible characters

65

Isaksson, M. and E. Mossel Maximally stable Gaussian partitions with discrete applications

347

Jain, V. K. and M. K. Yadav On finite p-groups whose automorphisms are all central

225

Kechris, A. S. Weak containment in the space of actions of a free group

461

Kim, C.-W. Finsler manifolds without conjugate points and with integral Ricci curvature

135

Krähmer, U. On the Hochschild (co)homology of quantum homogeneous spaces

237

Kühnel, W. See U. Brehm Lacroix, Y. See T. Downarowicz Leader, I. See B. Bollobás Marchese, A. and C. Zanco On a question by Corson about point-

finite coverings

55

McCutcheon, R. See P. Balister Mesyan, Z. Monoids of injective maps closed under conjugation by

permutations

287

Molina, S. Ribet bimodules and the specialization of Heegner points  1

Mossel, E. See M. Isaksson Navarro, G. See I. M. Isaacs

Peres, Y. and D. Ralston Heaviness in toral rotations  337

Raghavan, D. A model with no strongly separable almost disjoint families

39

Ralston, D. See Y. Peres Ruskuc, N. See R. Gray Sangroniz, J. See I. M. Isaacs

Shumyatsky, P. Multilinear commutators in residually finite groups  207

da Silva, P. R. See P. T. Cardin Teixeira, M. A. See P. T. Cardin Walters, M. See B. Bollobás Yadav, M. K. See V. K. Jain Zanco, C. See A. Marchese Zhang, S. Rigidity theorem for complete spacelike hypersurfaces with

constant scalar curvature in Lorentz spaces 177